OpenStudy (anonymous):

Which statement is true? Converting an improper fraction to a decimal shows if it is rational. Repeating decimals are never rational. A zero numerator over an integer denominator is never rational. A negative integer is always rational.

OpenStudy (anonymous):

last one..

OpenStudy (anonymous):

how do u know

OpenStudy (anonymous):

a. what's the definition of a rational number? b. How can rational numbers be expressed in decimal form? c. what's the definition of a rational number? d. what's the definition of a rational number?

OpenStudy (anonymous):

because integers are always rational whether it is - or +

OpenStudy (anonymous):

\[\mathbb{N} \subset \mathbb{Z} \subset \mathbb{Q} \subset \mathbb{R}\]

OpenStudy (anonymous):

@izzy_swaggman first understand the definitions..